OpenStudy (anonymous):

solve each equation. z^2 - 6z - 27 = 0 ?

OpenStudy (uri):

This is a Quadratic equation...right?

OpenStudy (anonymous):

@uri what have you been eating :)

OpenStudy (uri):

Nothing...it was ramadan remember...why? @dlearner

OpenStudy (anonymous):

\[z^2-9z-3z-27=0\]\[(z+3)(z-9)=0\]

OpenStudy (anonymous):

is that correct @deena123

OpenStudy (anonymous):

u can check it if u want @broderick365

OpenStudy (anonymous):

this was the wrong equation the equation is z^2 - 6z - 27 =0 ? @deena123

OpenStudy (anonymous):

sory a typo from my side it should be -9z+3z instead of -9z-3z

OpenStudy (uri):

z^2 - 6z - 27 The number when adding or subracting should be -6 so,-9+3=-6 and when multiplying should be 27 so,9 x 3=27 So, z^2-9z+3z-27 z(z-9)+3(z-9)

OpenStudy (uri):

Answer will be :(z-9)(z+3)